You purchased one share of Best Buy CO., Inc for $55.21 per share. The company paid a dividend of $9.24 per share during the year, and had an ending share price of $62. What is the percentage return?
Answer #1
Percentage return = (D+(P1-P0))/P0 Where,
= (9.24+(62.00-55.21))/55.21 D = $       9.24
= 29.03% P0 = $    55.21
P1 = $    62.00